# Ahmet Faruk Saz,  Siheng Xiong,  and Faramarz Fekri

### v1

@@ Abstract
This work presents an analysis of semantic communication in the context of First-Order Logic (FOL)-based deduction. Specifically, the receiver holds a set of hypotheses about the State of the World (SotW), while the transmitter has incomplete evidence about the true SotW but lacks access to the ground truth. The transmitter aims to communicate limited information to help the receiver identify the hypothesis most consistent with true SotW. We formulate the objective as approximating the posterior distribution at the transmitter to the receiver. Using Stirling’s approximation, this reduces to a constrained, finite-horizon resource allocation problem. Applying the Karush-Kuhn-Tucker conditions yields a truncated water-filling solution. Despite the problem’s non-convexity, symmetry and permutation invariance ensure global optimality. Based on this, we design message selection strategies, both for single- and multi- round communication, and model the receiver’s inference as an $m$-ary Bayesian hypothesis testing problem. Under the Maximum A Posteriori (MAP) rule, our communication strategy achieves optimal performance within budget constraints. We further analyze convergence rates and validate the theoretical findings through experiments, demonstrating reduced error over random selection and prior methods.
Index Terms:
First-order logic, semantic communication, water-filling, hypothesis testing, deduction
@@ I Introduction
Semantic communication redefines data transmission by prioritizing meaning over raw data, offering critical benefits for intelligent services and next-generation networks like 6G. Previous works have explored semantic compression techniques [1], knowledge-enhanced transmission strategies [2], and semantic-aware resource allocation. These efforts have shown the promise of semantic communication for improving efficiency and robustness across diverse intelligent systems and communication environments.
@@
While these works largely emphasize semantic compression and transmission optimization, our focus shifts to a distinct but complementary problem: logical deduction of hypotheses under resource constraints. In many real-world scenarios—such as distributed sensing or privacy-sensitive environments—the transmitter lacks knowledge of the receiver’s precise goal. Consequently, the challenge becomes transmitting the most semantically content-informative message possible. This framework has the potential to empower a variety of applications involving logical deduction: drones in search-and-rescue operations can share logical summaries of their local observations [3]; wearable health monitors can transmit structured symptom patterns to support remote diagnostics [4]; autonomous vehicles can communicate compact causal cues like “reduced traction ahead” [4]; and legal or educational systems can tailor logic-based explanations to user intent [5].
@@
This paper builds on prior work that introduced a First-Order Logic (FOL)-based semantic communication framework for deductive inference. Drawing from the foundational ideas of Carnap [6] and Hintikka [7], we treated FOL as a formal language for representing and transmitting structured knowledge. While classical semantic systems were computationally intractable, our earlier work [8, 9, 10] addressed scalability by framing communication as a boolean-satisfiability model selection problem. Using model counting, we developed a bit-efficient, task-agnostic transmission strategy that enabled logical inference at the receiver, even when the transmitter lacked access to the downstream task.
@@
In this paper, we take our efforts a step further and formalize semantic communication as a posterior distribution alignment problem over a logical space structured by First-Order Logic (FOL). The world is represented using a hierarchy of constituents, attributive constituents, and Q-sentences, and uncertainty is quantified via inductive logical probability over these structures. The transmitter, observing partial evidence about the true state, aims to select messages that maximize semantic informativeness—quantified via Carnap’s cont-information—without access to the receiver’s hypothesis set. We show that optimal message selection reduces to minimizing the KL divergence between transmitter and receiver posteriors over constituents, which in turn becomes a constrained resource allocation problem over observed attributive constituents. To make the analysis tractable, we apply Stirling’s approximation and exploit the symmetry of the resulting objective, deriving a globally optimal solution using a truncated water-filling algorithm. We extend this framework to multi-round communication, showing that repeated optimal allocations converge to an ideal posterior configuration under cumulative constraints. This formulation enables principled message selection under realistic communication budgets and, under the Bayesian MAP criterion at the receiver, minimizes error in hypothesis deduction. We derive the convergence rate for the multi-round setup and show that the algorithm converges linearly. Finally, we validate our theoretical findings on a custom deduction dataset, demonstrating superior performance in hypothesis selection compared to prior work, greedy selection, and random baselines.
@@
The contribution of this paper is twofold. First, it introduces a formal connection between semantic communication and logical inference by framing message transmission as the problem of posterior distribution alignment between transmitter and receiver in a First-Order Logic framework. This perspective distinguishes the present work from prior approaches that emphasize semantic compression or task-specific heuristics, and provides a principled way to quantify semantic informativeness through Carnap’s cont-information and distributional divergence. Second, we develop and analyze a resource allocation formulation of this problem, showing that the optimal message composition follows a truncated water-filling solution. We establish that this strategy achieves global optimality in the relaxed setting and extend it to multi-round communication, where convergence properties are derived. Simulation results confirm that aligning posterior distributions yields significant improvements in hypothesis deduction accuracy compared to prior semantic communication strategies. Together, these contributions advance semantic communication beyond compression-oriented formulations and provide a theoretically grounded approach for logic-based deductive reasoning under communication constraints. In the next section, we introduce FOL representation of the world and problem setup.
@@ II Related Work
Semantic and goal-oriented communication systems, inspired by Weaver’s Level B notion, focus on transmitting meaning rather than symbols. DeepSC and DeepJSCC [11, 12] pioneered autoencoder-based designs, while other works introduced task-oriented training [13, 14] and functional compression [15]. These methods achieve notable gains in efficiency but often rely on black training or static task definitions.
Beyond end-to-end learning, neuro-symbolic and graph-based approaches aim to combine logical reasoning with semantic transmission. Prior work [16, 17, 18, 19, 20] has shown that these methods strengthen semantic understanding by capturing structured meaning, making them promising foundations for communication systems. RuleTaker and ProofWriter [21, 22] demonstrated transformer-based reasoning, while recent studies integrate knowledge graphs and large models [23, 24] to encode structured semantics. Vector quantization [25] and generative models [26] further extend semantic communication by aligning semantic tokens with modulation schemes or reconstructing content via shared AI models.
@@ III-A Representing the State of the World via First-Order Logic
We begin by illustrating how the state of the world can be represented using First-Order Logic (FOL). Consider an infinite world $\mathbf{W}_{I}$ characterized by a countably infinite set of entities $\mathbf{E}$ and a finite set of predicates $\mathbf{P}$. Each FOL sentence $m_{i}$ is described using the logical language $\mathcal{L}$, which includes the predicates $\mathbf{P}$, entities $\mathbf{E}$, quantifiers $\forall,\exists$, and logical connectives $\land,\lor,\neg$. To describe the state in $\mathbf{W}_{I}$, we start by introducing the concept of Q-sentences.
@@
A Q-sentence $Q_{i}(a,b)$ is defined as:
$Q_{i}(a,b)=\bigwedge_{r\in\mathcal{I}_{P}}\delta_{r}R_{r}(a,b)\land\delta^{\prime}_{r}R_{r}(b,a),$ (1)
where $a,b\in\mathbf{E}$ and $\delta_{r},\delta^{\prime}_{r}\in\{+1,-1\}$ indicate the presence or negation of predicates.
Building upon Q-sentences, we define attributive constituents to capture the relationships between individual entities.
@@ A Q-sentence $Q_{i}(a,b)$ is defined as:
An attributive constituent $C_{z}(x)$ in $\mathbf{W}_{I}$ is:
$C_{z}(x)=\bigwedge_{i\in\mathcal{I}_{\mathcal{Q}}}\delta_{k}\{(Q_{i}(x,b)):\exists b\},$ (2)
where $\delta_{i}\in\{+1,-1\}$, $\mathcal{I}_{\mathcal{Q}}$ is the index set over Q-sentences, and $i=1,...|\mathcal{I}_{\mathcal{Q}}|$, and $b\in\mathbf{E}$. The expression encapsulates all relationships entity $x$ has within the world, expressing what kind of individual $x$ is.
To represent the state in an infinite world (in a generalized manner), we define constituents as combinations of attributive constituents.
@@
A constituent $C^{w}$ of width $w$ in $\mathbf{W}_{I}$ is a conjunction of attributive constituents:
$C^{w}=\bigwedge_{t=1}^{w}\;\{C_{z}(x):\exists x\},$ (3)
which represents the relational structure of entities in the world $\mathbf{W}_{I}$. Here, the width $w$ is the number of attributive constituents $C_{z}(x)$ present in the conjunction forming $C^{w}$.
Theorem 1 (FOL Representation in Infinite World [7]).
Any FOL statement $m$ in $\mathbf{W}_{I}$ can be represented as:
@@
where $\mathcal{I}_{\mathcal{C}}$ is the index set of constituents.
In FOL-based representation of world $\mathbf{W}_{I}$, we introduce a probability measure $\mathcal{P}_{I}$ over the fundamental elements, i.e., constituents $C^{j}$. To define this measure, we make use of inductive logical (a.k.a., epistemic) probabilities, which serve as inductive Bayesian posteriors, built upon observations and a prior distribution if any.
Definition 4 (Inductive Logical Probability [6]).
For any statement $g$ and evidence $e$ (observations) in $\mathbf{W}_{I}$, the inductive logical probability (a.k.a., degree of confirmation) $c(g,e)$ that evidence $e$ supports statement $g$ is defined as:
$c(g,e)=p(g\mid e)=\frac{p(g\land e)}{p(e)}=\frac{l_{g}+\lambda(w)/w}{l+\lambda(w)},$ (5)
where $l$ is the number of observations, $l_{g}$ is the number of observations confirming statement $g$, $w$ is the width of the constituent $C^{w}$ instantiated by statement $g$, $\lambda(w)\geq 0$ is the prior coefficient, possibly a function of $w$.
Then, the inductive logical probability measure $\mathcal{P}_{I}$ over $\mathcal{C}$ is defined as $\mathcal{P}_{I}=\{c(C^{j},e)\mid j\in\mathcal{I}_{\mathcal{C}}\}$ where $\sum_{j\in\mathcal{I}_{\mathcal{C}}}c(C^{j},e)=1$. To quantify the remaining uncertainty after new observations, Carnap proposed the concept of semantic ”cont(ent)-information”, defined as:
@@
The cont-information in a world $\mathbf{W}_{I}$ quantifies the informativeness of an observation $e$ to a specific observer by measuring the reduction in uncertainty about the SotW. Given a state $C^{j}$ and evidence $e$ (observations) in $\mathbf{W}_{I}$, cont-information is defined as:
$\text{cont}(C^{j};e)=1-c(C^{j},e)=1-p(C^{j}|e),$ (6)
In simple terms $\text{cont}(C^{j};e)$ measures the uncertainty reduction about constituent $C^{j}$ being the true SotW after observing e. Next, the cont-information measure is used to devise a semantic communication framework.

### v2

@@ Abstract
This work presents an analysis of semantic communication in the context of First-Order Logic (FOL)-based deduction. Specifically, the receiver holds a set of hypotheses about the State of the World (SotW), while the transmitter has incomplete evidence about the true SotW but lacks access to the ground truth. The transmitter aims to communicate limited information to help the receiver identify the hypothesis most consistent with true SotW. We formulate the objective as approximating the posterior distribution of the transmitter at the receiver. Using Stirling’s approximation, this reduces to a constrained, finite-horizon resource allocation problem. Applying the Karush-Kuhn-Tucker conditions yields a truncated water-filling solution. Despite the problem’s non-convexity, symmetry and permutation invariance ensure global optimality. Based on this, we design message selection strategies, both for single- and multi- round communication, and model the receiver’s inference as an $m$-ary Bayesian hypothesis testing problem. Under the Maximum A Posteriori (MAP) rule, our communication strategy achieves optimal performance within budget constraints. We further analyze convergence rates and validate the theoretical findings through experiments, demonstrating reduced error over random selection and prior methods.
@@ I Introduction
Semantic communication redefines data transmission by prioritizing meaning over raw data, offering critical benefits for intelligent services and next-generation networks like 6G. Previous works have explored semantic compression techniques [Gund_Yen], knowledge-enhanced transmission strategies [Qin2021SemanticCP], and semantic-aware resource allocation. These efforts have shown the promise of semantic communication for improving efficiency and robustness across diverse intelligent systems and communication environments.
@@
While these works largely emphasize semantic compression and transmission optimization, our focus shifts to a distinct but complementary problem: logical deduction of hypotheses under resource constraints. In many real-world scenarios—such as distributed sensing or privacy-sensitive environments—the transmitter lacks knowledge of the receiver’s precise goal. Consequently, the challenge becomes transmitting the most semantically content-informative message possible. This framework has the potential to empower a variety of applications involving logical deduction: drones in search-and-rescue operations can share logical summaries of their local observations [Javaid2024LargeLM]; wearable health monitors can transmit structured symptom patterns to support remote diagnostics [Mahmud]; autonomous vehicles can communicate compact causal cues like “reduced traction ahead” [Mahmud]; and legal or educational systems can tailor logic-based explanations to user intent [Singamaneni2023ASO].
@@
This paper builds on prior work that introduced a First-Order Logic (FOL)-based semantic communication framework for deductive inference. Drawing from the foundational ideas of Carnap [c2] and Hintikka [c10], we treated FOL as a formal language for representing and transmitting structured knowledge. While classical semantic systems were computationally intractable, our earlier work [saz2024lossy, saz2024model, saz2025discd] addressed scalability by framing communication as a boolean-satisfiability model selection problem. Using model counting, we developed a bit-efficient, task-agnostic transmission strategy that enabled logical inference at the receiver, even when the transmitter lacked access to the downstream task.
@@
In this paper, we take our efforts a step further and formalize semantic communication as a posterior distribution alignment problem over a logical space structured by First-Order Logic (FOL). The world is represented using a hierarchy of constituents, attributive constituents, and Q-sentences, and uncertainty is quantified via inductive logical probability over these structures. The transmitter, observing partial evidence about the true state, aims to select messages that maximize semantic informativeness without access to the receiver’s hypothesis set. We show that optimal message selection reduces to minimizing the KL divergence between transmitter and receiver posteriors over constituents, which in turn becomes a constrained resource allocation problem over observed attributive constituents. To make the analysis tractable, we apply Stirling’s approximation and exploit the symmetry of the resulting objective, deriving a globally optimal solution using a truncated water-filling algorithm. We extend this framework to multi-round communication, showing that repeated optimal allocations converge to an ideal posterior configuration under cumulative constraints. This formulation enables principled message selection under realistic communication budgets and, under the Bayesian MAP criterion at the receiver, minimizes error in hypothesis deduction. We derive the convergence rate for the multi-round setup and show that the algorithm converges linearly. Finally, we validate our theoretical findings on a custom deduction dataset, demonstrating superior performance in hypothesis selection compared to prior work, greedy selection, and random baselines.
@@
The contribution of this paper is twofold. First, it introduces a formal connection between semantic communication and logical inference by framing message transmission as the problem of posterior distribution alignment between transmitter and receiver in a First-Order Logic framework. This perspective distinguishes the present work from prior approaches that emphasize semantic compression or task-specific heuristics, and provides a principled way to quantify semantic informativeness through distributional divergence. Second, we develop and analyze a resource allocation formulation of this problem, showing that the optimal message composition follows a truncated water-filling solution. We establish that this strategy achieves global optimality in the relaxed setting and extend it to multi-round communication, where convergence properties are derived. Simulation results confirm that aligning posterior distributions yields significant improvements in hypothesis deduction accuracy compared to prior semantic communication strategies. Together, these contributions advance semantic communication beyond compression-oriented formulations and provide a theoretically grounded approach for logic-based deductive reasoning under communication constraints. In the next section, we introduce FOL representation of the world and problem setup.
@@ II Related Work
Semantic and goal-oriented communication systems, inspired by Weaver’s Level B notion, focus on transmitting meaning rather than symbols. DeepSC and DeepJSCC [xie2021deep, bourtsoulatze2019deep] pioneered autoencoder-based designs, while other works introduced task-oriented training [shao2022task, yang2023task] and functional compression [Saidutta2022AML]. These methods achieve notable gains in efficiency but often rely on black-box training or static task definitions.
Beyond end-to-end learning, neuro-symbolic and graph-based approaches aim to combine logical reasoning with semantic transmission. Prior work [xiongtilp, yang2024temporal, xiong2024teilp, xiong2024large, xiong2024deliberate] has shown that these methods strengthen semantic understanding by capturing structured meaning, making them promising foundations for communication systems. RuleTaker and ProofWriter [clark2020transformers, tafjord2021proofwriter] demonstrated transformer-based reasoning, while recent studies integrate knowledge graphs and large models [guo2023kg, nour2024semantic] to encode structured semantics. Vector quantization [fu2023vector] and generative models [grassucci2023generative] further extend semantic communication by aligning semantic tokens with modulation schemes or reconstructing content via shared AI models.
@@ III-A Representing the State of the World via First-Order Logic
We begin by illustrating how the state of the world can be represented using First-Order Logic (FOL) following the conventions presented in [c4, nelte]. Consider a language $\mathcal{L}$ whose syntax is defined over vocabulary consisting of a finite set of dyadic predicate symbols $\mathbf{P}=\{P_{i}\}_{i=1}^{k}$, a countably infinite set of variables $\mathbf{X}=\{x_{i}\}_{i\in\mathbb{N}}$, a countable set of constant symbols $\mathbf{C}=\{\top,\bot,$...\}, the logical connectives $\{\neg,\land,\lor\}$, the quantifiers $\{\forall,\exists\}$, and parentheses $()$ for disambiguation.
The syntax of a formal language has no inherent meaning on its own; meaning is provided through a model. A model for a language ($\mathcal{L}$) is a pair $\mathcal{M}=\langle\mathbf{E},J\rangle$, where $\mathbf{E}$ is a nonempty, countably infinite set called the domain (or universe) whose elements are referred to as entities, and $J$ is an interpretation function that assigns constant symbols to entities in $\mathbf{E}$ and each binary predicate symbol $P\in\mathbf{P}$ a binary relation $J(P)\subseteq\mathbf{E}\times\mathbf{E}$.
To evaluate formulas that contain variables, a valuation function $v$ assigns domain elements to variables $v[x_{i}\mapsto e_{i}]$. Given a model and a valuation, the satisfaction relation ($\mathcal{M},v\models F$) formally specifies whether a formula ($F$) is true in the model $\mathcal{M}$ under the corresponding valuation $v$. The state of the world is then represented through true but unknown model-valuation pair $\mathcal{M}^{*},v^{*}$, which exhaustively determines the structure of the domain $\mathbf{E}$ and the complete pattern of relationships among its elements, thereby logically entailing all correct statements of $\mathcal{L}$.
As an initial step towards characterizing the state of the world in the framework explained, we start by defining specific predicate-relational configurations between entities through Q-sentences.
@@
For a fixed configuration $i$, let $\delta_{i,P},\,\delta^{\prime}_{i,P}\in\{+1,-1\},\forall\,P\in\mathbf{P}$. For $x_{1},x_{2}\in\mathbf{X}$, define the Q-sentence $Q_{i}(x_{1},x_{2})$ by
$Q_{i}(x_{1},x_{2})=\bigwedge_{P\in\mathbf{P}}\left(\delta_{i,P}P(x_{1},x_{2})\land\delta^{\prime}_{i,P}P(x_{2},x_{1})\right).$
Remark 1.
The coefficients $\delta_{i,P},\delta^{\prime}_{i,P}\in\{+1,-1\}$ encode whether the predicate $P$ appears positively or negatively in the Q-sentence. Concretely, $\delta_{i,P}=+1$ corresponds to the atomic formula $P(x_{1},x_{2})$, while $\delta_{i,P}=-1$ corresponds to its negation $\neg P(x_{1},x_{2})$; analogously, $\delta^{\prime}_{i,P}$ governs the occurrence of $P(x_{2},x_{1})$. Thus, each configuration $i$ specifies a complete choice of affirmation or negation for every predicate in both argument orders.
This definition implies that there are $4^{|\mathbf{P}|}$ distinct Q-sentences. For a fixed valuation $v[x_{1}\mapsto e_{1},x_{2}\mapsto e_{2}]$, exactly one of these $4^{|\mathbf{P}|}$ Q-sentences evaluates to 1 (true), while all others evaluate to 0 (false). Building on this observation, we introduce the notion of an attributive constituent, which captures the complete relational profile of a given entity.
@@
Let $\mathbf{Q}$ be the set of all Q-sentences. For a fixed configuration $j$, let $\delta_{j,Q}\in\{+1,-1\},\forall\,Q\in\mathbf{Q}$. For $x_{1}\in\mathbf{X}$, an attributive constituent $C_{j}(x_{1})$ is defined as:
$C_{j}(x_{1})=\bigwedge_{Q\in\mathbf{Q}}\delta_{j,Q}\bigl((\exists x_{2})\,Q(x_{1},x_{2})\bigr)$ (1)
Remark 2.
As before, the coefficients $\delta_{j,Q}\in\{+1,-1\}$ encode polarity: $\delta_{j,Q}=+1$ denotes the formula $(\exists x_{2})Q(x_{1},x_{2})$, while $\delta_{j,Q}=-1$ denotes its negation $\neg(\exists x_{2})Q(x_{1},x_{2})$. $\neg(\exists x_{2})Q(x_{1},x_{2})$ asserts that there exists no $x_{2}$ for which $Q(x_{1},x_{2})$ evaluates to true.
This expression encapsulates the complete set of relationships that an entity $x_{1}$ holds within the world, thereby characterizing what kind of individual $x_{1}$ is with respect to the relationships it possesses or lacks with others. Specifically, $C_{j}(x_{1})$ specifies the pattern of $Q$-relations that the individual $x_{1}$ has with all other entities in the universe. It therefore explicitly identifies which Q-sentences, among the $4^{|\mathbf{P}|}$ possible ones, are satisfied by $x_{1}$, as well as which are not.
For every individual $x_{1}$ in the universe, there exists a corresponding attributive constituent $C_{j}(x_{1})$. While distinct individuals may share the same attributive constituent, the total number of distinct attributive constituents is bounded above by $2^{4^{|\mathbf{P}|}}$ by definition.
To represent the state of the world in $\mathbf{E}$, we define constituents as combinations of attributive constituents.
@@
Let $\mathbf{C}_{\mathbf{A}}$ be the set of all attributive constituents. For a fixed configuration $k$, let $\delta_{k,C_{A}}\in\{+1,-1\},\quad\forall\,C_{A}\in\mathbf{C}_{\mathbf{A}}$. A constituent $C^{w}$ of width $w$ in $\mathcal{C}$ is a conjunction of attributive constituents:
$C^{w,k}=\bigwedge_{C_{A}\in\mathbf{C}_{\mathbf{A}}}\delta_{k,C_{A}}\bigl((\exists x_{1})C_{A}(x_{1})\bigr),$ (2)
The width $w$ is the number of attributive constituents in the conjunction with positive polarity $\delta_{k,C_{A}}=+1$.
Remark 3.
As before, the coefficients $\delta_{k,C_{A}}\in\{+1,-1\}$ encode polarity: $\delta_{k,C_{A}}=+1$ denotes the formula $(\exists x_{1})C_{A}(x_{1})$, while $\delta_{k,C_{A}}=-1$ denotes its negation $\neg(\exists x_{1})C_{A}(x_{1})$. $\neg(\exists x_{1})\,C_{A}(x_{1})$ asserts that there exists no $x_{1}$ for which $C_{A}(x_{1})$ evaluates to true.
This definition implies that there are at most $2^{2^{4^{|\mathbf{P}|}}}$ distinct constituents. The constituents $C^{w,k}$, for $k=1,\ldots$, enumerate all normal forms that can be constructed from the primitive predicate set $\mathbf{P}$, propositional connectives, and at most two quantifiers. Each constituent explicitly specifies which attributive constituents in $\mathbf{C}_{\mathbf{A}}$ are realized in the universe, thereby fully characterizing the relationship patterns that exist in $\mathbf{E}$.
Although up to $2^{2^{4^{|\mathbf{P}|}}}$ constituents can be defined for a universe $\mathbf{E}$, exactly one of them is satisfied. Let $\mathcal{C}$ be the set of all constituents that can be defined in $\mathcal{L}$ for the universe $\mathbf{E}$. For any true model–valuation pair $(\mathcal{M}^{*},v^{*})$ over $\mathbf{E}$, there exists a unique constituent $C^{*}\in\mathcal{C}$ such that $\mathcal{M}^{*}\models C^{*}$. This constituent represents the true state of the world.
We therefore interpret constituents as potential candidates for the state of the world, with the satisfied constituent corresponding to the actual state. In practice, determining the true state of the world thus amounts to identifying the constituent that is most consistent—i.e., best supported—by the available evidence $\mathcal{E}$.
Remark 4.
Multiple notions of the state of the world are possible, and the definition adopted above represents only one such formulation. In this work, we define the state of the world exclusively in terms of the relationship patterns that exist in $\mathbf{E}$. Alternative formulations are possible; for example, one may define the state of the world by characterizing, for each individual $x_{1}$, the type of individual $x_{1}$ is with respect to the relationships it holds with others (i.e., as the set of attributive constituents $\mathbf{C}_{\mathbf{A}}$ that holds true), or by adopting hybrid definitions that combine both perspectives, as in [c4].
Every general sentence (and hence every general hypothesis) $\varphi$ of $\mathcal{L}$ over $\mathbf{E}$ can be expressed as a disjunction of constituents, as established in the theorem below.
Theorem 1 (Distributive Normal Forms of General Sentences in FOL [nelte, c9, c4]).
Let $\mathcal{L}$ be a first-order language with a finite set of dyadic predicate symbols $\mathbf{P}$, and let every model $\mathcal{M}=\langle\mathbf{E},J\rangle$ for $\mathcal{L}$ have a countably infinite domain $\mathbf{E}$. Let $\mathcal{C}$ be the set of constituents described in $\mathcal{L}$. For any closed first-order general sentence $\varphi\in\mathcal{L}$, consider the set
$\{\,C^{w,k}\in\mathcal{C}\;:\;C^{w,k}\models\varphi\,\}$
such that
$\varphi\;\equiv\;\bigvee_{\begin{subarray}{c}C^{w,k}\in\mathcal{C}\\ C^{w,k}\models\varphi\end{subarray}}C^{w,k}.$ (3)
@@
In the first-order logic (FOL)–based representation of the world $\mathbf{E}$, we introduce a probability measure $\mathcal{P}_{I}$ over the constituents $C^{w,k}$. This measure is defined using inductive logical (also known as epistemic) probabilities, which function as inductive Bayesian posteriors constructed from observed evidence and, when available, a prior distribution.
Definition 4 (Inductive Logical Probability [c2]).
For any statement $\varphi$ and evidence $e$ (observations) in $\mathbf{E}$, the inductive logical probability (also referred to as the degree of confirmation) $c(\varphi,e)$ of $\varphi$ given $e$ is defined as
$c(\varphi,e)\;=\;p(\varphi\mid e)\;=\;\frac{p(\varphi\land e)}{p(e)}.$ (4)
We then define an inductive logical probability measure $\mathcal{P}_{I}$ over the set of constituents $\mathcal{C}$ as
$\mathcal{P}_{I}\;=\;\bigl\{\,c\!\left(C^{w,k},e\right)\;\big|\;C^{w,k}\in\mathcal{C}\,\bigr\},$
where $c(C^{w,k},e)$ denotes the inductive logical probability of the constituent $C^{w,k}$ given the evidence $e$, and
$\sum_{C^{w,k}\in\mathcal{C}}c\!\left(C^{w,k},e\right)\;=\;1.$
Once the probabilities of the constituents are fixed, the probability—and hence the inductively determined truth—of any hypothesis can be computed in accordance with Theorem (1).
It is important to note that the assignment of probabilities to constituents is relative to the inductive system adopted. Different systems—such as Carnap’s $c$-system [carnap_logical_1962], the $\lambda$-continuum [c1], and Hintikka’s two-dimensional framework [c10]—employ different ways of determining the probability assignments [c1]. Consequently, these systems induce different probability distributions over constituents. In the present work, we adopt the conventions and framework developed in [c4].
To quantify the remaining uncertainty after incorporating new observations, Carnap introduced the notion of semantic content-information, defined as follows:
@@
The cont-information of a sentence quantifies the informativeness of an observation $e$ for a given observer in a world $\mathbf{E}$ by measuring the reduction in uncertainty about a general sentence (i.e., a general hypothesis) $\varphi$ of the language $\mathcal{L}$. Given a sentence $\varphi\in\mathcal{L}$ and evidence $e$ (observations) in $\mathbf{E}$, the cont-information of $\varphi$ given $e$ is defined as
$\text{cont}(\varphi;e)\;\coloneqq\;1-c(\varphi,e)\;=\;1-p(\varphi\mid e).$ (5)
In intuitive terms, $\text{cont}(\varphi;e)$ measures the reduction in uncertainty about the truth of a sentence $\varphi$ after observing the evidence $e$. In the following, the cont-information measure is used to develop a semantic communication framework.